OpenStudy (anonymous):

I need to be walked, step by step to answer this question, Please! angle 1 & angle 2 are supplementary. angle 1 is 4y+7 and angle 2 is 9y+4. find angle 2

OpenStudy (radar):

What do supplementary angles add up to...........How many degrees?

OpenStudy (anonymous):

180

OpenStudy (radar):

Yes, I agree. So we know that angle 1 + angle 2 = 180 degrees. Agree?

OpenStudy (anonymous):

yes

OpenStudy (radar):

We now got it made. Angle 1 is 4y + 7 and angle 2 is 9y + 4 and their sum is 180. We can now write: (4y = 7) + (9y + 4) = 180. You can combine similar terms, simplify and solve for y. After getting y value plug that in for angle 2 and solve for the angle in degrees.

OpenStudy (radar):

Do you follow that?

OpenStudy (anonymous):

136 degrees???

OpenStudy (radar):

Gee that was quick. Just let me check. 13y + 11 = 180 13 y = 180 - 11 = 169 y = 13 9y + 4 = 9(13) + 4 = 117 + 4 = 121 degree That is what I got for angle 2. Let me check by finding angle 1 and add the angles and see if they add up to 180 4 (13) + 7 = 52 + 7 = 59 Adding the two 121 + 59 = 180. Check your solution for y.

OpenStudy (radar):

What did you get for y?

OpenStudy (anonymous):

14.692, & thank you so much!

OpenStudy (anonymous):

I see. I added the 11, instead of subtracting it!

OpenStudy (anonymous):

thank you & have a blessed day :)

OpenStudy (radar):

No No, not 14.692. Here is it step by step